OpenStudy (anonymous):

Change the exponential equation to an equivalent equation involving a logarithm 2401=7^4

OpenStudy (anonymous):

\[b^k = a \iff log_ba = k\] The arrows mean you can go back and forth from one to the other.

OpenStudy (anonymous):

In your example here you have: \[7^4 = 2401 \iff ?\]

OpenStudy (anonymous):

i realize they are equal because 7^4=2401. I am not sure how to write it in logarithmic form

OpenStudy (anonymous):

I'm not saying anything about their equality.

OpenStudy (anonymous):

Look at what I wrote above. It explains how to go back and forth from exponential to logarithmic forms: \[b^k = a \iff log_ba = k\]

OpenStudy (anonymous):

The arrows indicate that you can alternate between one form and the other. They both mean the same thing.

OpenStudy (anonymous):

Perhaps it's hard to make out that those are arrows.. I'll make em bigger: \[\huge b^k = a \iff log_ba =k \]

OpenStudy (anonymous):

In your case 'b' is 7, 'k' is 4, and 'a' is 2401

OpenStudy (anonymous):

Did you figure it out?

OpenStudy (anonymous):

If the letters confuse you I can use numeric examples.

OpenStudy (anonymous):

Yes i understood all that thank you very much. I just get confused on final answers. i do my homework online so if i make mistakes it wont give me credit

OpenStudy (anonymous):

if i put in the wrong form

OpenStudy (anonymous):

I put log[7]^2401=4 as final answer

OpenStudy (anonymous):

That's correct. \[log_7(2401) = 4\]

OpenStudy (anonymous):

Thank you =)